# CasualHDRSplat: Robust High Dynamic Range 3D Gaussian Splatting from Casually Captured Videos

### v1

@@ Abstract
Recently, photo-realistic novel view synthesis from multi-view images, such as neural radiance field (NeRF) and 3D Gaussian Splatting (3DGS), have garnered widespread attention due to their superior performance. However, most works rely on low dynamic range (LDR) images, which limits the capturing of richer scene details. Some prior works have focused on high dynamic range (HDR) scene reconstruction, typically require capturing of multi-view sharp images with different exposure times at fixed camera positions during exposure times, which is time-consuming and challenging in practice. For a more flexible data acquisition, we propose a one-stage method: CasualHDR to easily and robustly reconstruct the 3D HDR scene from casually captured videos with auto-exposure enabled, even in the presence of severe motion blur and varying unknown exposure time. CasualHDR contains a unified differentiable physical imaging model which first applies continuous-time trajectory constraint to imaging process so that we can jointly optimize exposure time, camera response function (CRF), camera poses, and sharp 3D HDR scene. Extensive experiments demonstrate that our approach outperforms existing methods in terms of robustness and rendering quality.
https://github.com/WU-CVGL/CasualHDRSplat
@@
Figure 1: a) Our method can reconstruct 3D HDR scenes from videos casually captured with auto-exposure enabled. b) Our approach achieves superior rendering quality compared to methods like Gaussian-W and HDR-Plenoxels. c) After 3D HDR reconstruction, we can not only synthesize novel view, but also perform various downstream tasks, such as 1) HDR exposure editing, 2) Image deblurring.
11footnotetext: Equal contribution.
1 Introduction
@@
Photo-realistic Novel View Synthesis (NVS) are vital areas in computer vision with applications in VR/AR, autonomous driving, and embodied AI, offering immersive experiences for both humans and AI agents. Neural Radiance Fields (NeRFs)[36] and 3D Gaussian Splatting (3DGS)[18], through reconstructing radiance fields, have become mainstream methods in the NVS field due to their high-quality rendering, inspiring a large number of subsequent research works[40, 64, 54, 13].
However, most NVS methods struggle with high-contrast inputs, assuming good-quality images with consistent exposure conditions and low dynamic range (LDR). On the other hand, limited dynamic range of the inputs hinders 3D scene representations from reconstructing fine details in high dynamic range (HDR) environments, thereby restricting its further applications. Although 2D HDR contents have been standardized, consumed and exploited in recent years [12, 15, 1], 3D HDR free-viewpoint (volumetric) content is still a new concept with great potential. Therefore, reconstructing 3D high dynamic range (HDR) scenes is of significant practical value for achieving better visual effects and meeting the needs of downstream tasks.
@@
Current 3D HDR NVS methods can be divided into two categories. The first category, e.g. RawNeRF[37] and LE3D[16] etc., takes in noisy RAW images, aiming to reconstruct noise-free 3D HDR scenes. The second category, represented by [13, 4, 17, 53, 58], draws inspiration from HDR imaging (HDRI), using multi-exposure LDR images as inputs to reconstruct the 3D HDR scene while learning camera response function (CRF). However, the strict inputs limit their flexibility and broader applications. The challenges include: 1) Data acquisition of RAW images and accurate exposure time is usually expensive due to the use of professional equipment; 2) In low-light conditions, long exposure times increase the risk of motion blur from camera shake, reducing reconstruction quality; 3) The geometric consistency will be compromised if given inaccurate camera pose initialization, as the camera poses are not being optimized. Thus, a key challenge is reducing the cost of data acquisition, enabling high-quality and robust 3D HDR scene reconstruction with consumer-grade devices.
@@
Most modern consumer-grade cameras use auto-exposure during video recording, automatically adjusting exposure time based on ambient lighting. This expands the captured dynamic range in the video, making it possible to reconstruct 3D HDR scenes. However, naively applying these videos to existing HDR 3D reconstruction methods presents several challenges: 1) Although some cameras support providing exposure time in images stored in EXIF metadata, it is often impossible to obtain the exposure time for each frame in a video format; 2) Auto-exposure can cause inconsistencies in brightness between frames, leading to pose estimation errors in structure from motion (SfM) frameworks; 3) Videos suffer from severe motion blur due to the camera movement during exposure time.
To address these challenges, we cannot assume that the camera is static during the exposure time, as done in previous methods. Instead, we must also account for camera motion during this period. By analyzing the physical imaging process, we found that motion blur and image brightness variations are both directly related to the exposure time: a longer exposure time typically leads to more severe motion blur and higher image brightness. In previous works [13, 54], it has been shown that camera motion blur and image brightness can be decoupled as camera movement and the accumulation of irradiance during the exposure time. Thus, camera motion blur and varying image brightness can be constrained by the exposure time.
Building on above reasoning, we first introduced the continuous-time trajectory to represent the camera motion during video recording rather than addressing each camera pose as previous works[54, 64], so that the camera poses and exposure times can be jointly optimized. Subsequently, we design a camera response function (CRF) module which can convert the accumulation of irradiance (HDR image) over the exposure time to image brightness (LDR image). As mentioned above, exposure time links motion blur and image brightness. We incorporate camera motion into the physical imaging model and build a unified differentiable rendering framework, allowing the trajectory, CRF, and exposure time to be jointly optimized and constrained.
With explorations above, we propose a one-stage method called CasualHDRSplat, a unified 3D HDR reconstruction framework that couples the physical imaging model with camera motion, improving the robustness and flexibility. In our designed unified imaging model, the continuous-time camera trajectory on $\mathbb{SE}(3)$, exposure time, and camera response function are jointly optimized and mutually constrained. Therefore, our approach doesn’t require ground truth exposure times as inputs like previous methods.
@@
  • CasualHDRSplat, a unified imaging model that first applies continuous-time trajectory to constrain physical imaging model so that camera poses, CRF, exposure times and 3D HDR representation can be jointly optimizes, enabling users to reconstruct 3D HDR scenes from casually captured videos at a low cost.
  • A dataset that includes both synthetic data and real-world data, where each video contains severe variations in brightness and camera motion blur, that can be useful to the community to further investigate into this problem.
@@
  • With extensive experiments, we demonstrate how to utilize this model to reconstruct high-quality HDR scenes from casual videos, and exhibit state-of-the-art performance across all datasets.
2 Related Work
2.1 High Dynamic Range Imaging
@@
High Dynamic Range Imaging (HDRI) enhances luminosity beyond standard digital imaging by merging multi-exposure LDR images. In video capture, alternating long and short exposures achieves similar effects. Recently, deep learning approaches have treated HDRI as an image domian translation task, designing networks to convert LDR to HDR images. However, camera motion during exposure time often lead to ghosting artifacts. To address this, [11] proposed an adaptive metering algorithm to adjust exposure and reduce motion artifacts, while other methods use spatial attention to mitigate motion blur. With the advent of 3D scene representations such as NeRF and 3DGS, methods like [13, 17, 4, 14, 53] have emerged to reconstruct 3D HDR scenes and calibrate CRFs simultaneously. While these methods are effective, they often rely on precise exposure times and struggle with motion blur, highlighting the need for improved robustness.
2.2 Image deblurring
@@
Image deblurring aims to restore sharp images from blurred ones and can be categorized into three types. The first type uses hand-crafted priors, like total variation and heavy-tailed gradient priors, to solve for the blur kernel [21, 6], but struggles with different kernels producing similar effects. The second type, deep learning-based, achieves end-to-end restoration using large datasets [61, 51]. The third type leverages multi-view blurred images to reconstruct the 3D scene and deblur the images with geometric constraints. Works like [34, 42, 22] jointly learn the blur kernel and radiance field, while BAD-NeRF[54] introduces a physical motion blur model that jointly optimizes the radiance field and camera trajectories. Following BAD-NeRF[54], emerging works [23, 26, 25, 24, 49, 5, 41, 64, 60, 27, 43, 50] have demonstrated the effectiveness of using $\mathbb{SE}(3)$ trajectory representations for joint 3D reconstruction and blur image recovery.
2.3 Robust Novel View Synthesis
@@
Novel view synthesis aims to generate images from arbitrary viewpoints using input images with known poses. Neural Radiance Fields (NeRF)[35] and 3D Gaussian Splatting (3DGS)[18] advanced this field by reconstructing 3D scene as radiance field and achieve great success.
Most synthesis methods assume high-quality input data, but quality degrades significantly with blurry images, large exposure variations, or inaccurate poses. NeRF-W[35] and Gaussian-W[62] address these challenges by adding optimizable appearance vectors to model varied conditions. HDR-NeRF[13] and HDR-Plenoxels[17] reconstruct HDR scenes even with dynamic exposure variations. Other methods, like [8], handle pose inaccuracies, while BAD-Gaussians[64] incorporates camera motion models to deblur inputs during reconstruction.$I^{2}$-SLAM[2] is a concurrent approach capable of processing images with exposure inconsistencies and blur, though it focuses on RGB-D SLAM and has different representations of trajectory and CRF. These methods collectively improve robustness for NVS under challenging conditions.
3 Method
@@
In this section, we will provide a detailed explanation of our proposed CasualHDRSplat, which takes video captured with auto-exposure enabled as input. We first use continuous-time trajectory to represent the camera motion during video recording rather than addressing each camera pose individually, so that the camera pose and exposure time can be jointly optimized through bundle adjustment. Subsequently, we design a camera response function (CRF) module which can convert the accumulation of irradiance (HDR image) over the exposure time to pixel value (LDR image). Due to the fact that the exposure time is the connection between the continuous-time trajectory and CRF module, we can treat them as a unified differentiable imaging model thus the trajectory, CRF and exposure time can be jointly optimized and constrained together.
3.1 Preliminary: 3D Gaussian Splatting
@@
$\mathbf{G}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{\top}\mathbf{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})}$ (1)
$\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T},\mathbf{\Sigma^{\prime}}=\mathbf{J}\mathbf{R}_{c}\mathbf{\Sigma}\mathbf{R}_{c}^{T}\mathbf{J}^{T},$ (2)
@@
$\mathbf{C}(x,y,\mathbf{P}_{i})=\sum_{i=1}^{N}\mathbf{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (3)
$\alpha_{i}=\mathbf{o}_{i}\cdot\exp(-\sigma_{i}),\quad\sigma_{i}=\frac{1}{2}\Delta_{i}^{T}\mathbf{\Sigma^{\prime}}^{-1}\Delta_{i},$ (4)
@@
Figure 2: The pipeline of CasualHDRSplat. Given a casually captured video with auto exposure, camera motion blur, and significant exposure time changes, we train 3DGS to reconstruct an HDR scene. We design a unified model based on the physical image formation process, integrating camera motion blur and exposure-induced brightness variations. This allows for the joint estimation of camera motion, exposure time, and camera response curve while reconstructing the HDR scene. After training, our method can sharpen the train images and render HDR and LDR images from specified poses.
3.2 Continuous Trajectory Representation
@@
Cumulative $\mathbb{SE}(3)$ B-spline is a widely used continuous-time trajectory representation in robotics, especially in state estimation, sensor fusion and path planning [9, 33, 3, 45, 39, 10]. $\mathbb{SE}(3)$ B-spline has many excellent characteristics: such as $C^{2}$ continuity, locality and convex hull property that delicately incorporates gradient information and dynamic constraints, which converges quickly to generate smooth and feasible trajectories[65]. $\mathbb{SE}(3)$ B-spline allows for the calculation of pose, velocity, and accelerations at any timestamp given along a trajectory.
@@
Targeting at unordered inputs, existing multi-view deblurring methods following [54] model the camera motion and estimate short splines for each frame separately, thus cannot utilize the cross-frame motion constrains and priors, given a continuous video as their input. Some methods [52, 28, 49, 29, 30, 47, 55] utilize basis functions to regularize continuous-time deformations to reconstruct dynamic scenes, but robust reconstruction from casual videos with a continuous-time camera trajectory representation has not yet been explored. To this end, this paper estimates the camera motion across the whole video with a continuous-time cumulative $\mathbb{SE}(3)$ B-spline trajectory.
@@
Following [33], given a series of temporally uniformly distributed control knots, the pose $\mathbf{P}(t)$ at a given timestamp $t$ can be interpolated with 4 adjacent control knots, denoted as $\mathbf{T}_{\mathrm{0}}$, $\mathbf{T}_{1}$, $\mathbf{T}_{2}$ and $\mathbf{T}_{3}\in\mathbb{SE}(3)$:
$\mathbf{P}(t)=\mathbf{T}_{0}\cdot\prod_{j=0}^{2}\mathrm{exp}(\tilde{\mathbf{B}}(u)_{j+1}\cdot\bm{\Omega}_{j}),$ (5)
$\tilde{\mathbf{B}}(u)=\mathbf{C}\begin{bmatrix}1\\ u\\ u^{2}\\ u^{3}\end{bmatrix},\quad\mathbf{C}=\frac{1}{6}\begin{bmatrix}6&0&0&0\\ 5&3&-3&1\\ 1&3&3&-2\\ 0&0&0&1\end{bmatrix}.$ (6)
@@
where $\tau$ represents the spline sampling interval, $u=\frac{t}{\tau}$, and $u$ lies within the interval $[0,1)$; $\tilde{\mathbf{B}}(u)_{j+1}$ denotes the $(j+1)^{th}$ element of the vector $\tilde{\mathbf{B}}(u)$, $\bm{\Omega}_{j}=\mathrm{log}(\mathbf{T}_{j}^{-1}\cdot\mathbf{T}_{j+1})$, based on the [44].
3.3 Physical Image Formation Model

### v2

# Casual3DHDR: High Dynamic Range 3D Gaussian Splatting from casually-captured Videos

@@ Abstract.
Photo-realistic novel view synthesis from multi-view images, such as neural radiance field (NeRF) and 3D Gaussian Splatting (3DGS), has gained significant attention for its superior performance. However, most existing methods rely on low dynamic range (LDR) images, limiting their ability to capture detailed scenes in high-contrast environments. While some prior works address high dynamic range (HDR) scene reconstruction, they typically require multi-view sharp images with varying exposure times captured at fixed camera positions—a process that is time-consuming and impractical. To make data acquisition more flexible, we propose Casual3DHDR, a robust one-stage method that reconstructs 3D HDR scenes from casually-captured auto-exposure (AE) videos, even under severe motion blur and unknown, varying exposure times. Our approach integrates a continuous camera trajectory into a unified physical imaging model, jointly optimizing exposure times, camera poses, and the camera response function (CRF). Extensive experiments on synthetic and real-world datasets demonstrate that Casual3DHDR outperforms existing methods in robustness and rendering quality.
High dynamic range, Motion blur, 3D Gaussian Splatting, Novel view synthesis, Casual video, Exposure time, Camera response function, Pose optimization, Camera trajectory, B-Spline curve
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755612isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Reconstructionccs: Computing methodologies Image-based renderingccs: Computing methodologies 3D imaging
@@
Figure 1. a) Our method can reconstruct 3D HDR scenes from videos casually-captured with AE enabled. b) Our approach achieves superior rendering quality compared to methods like Gaussian-W and HDR-Plenoxels. c) After 3D HDR reconstruction, we can not only synthesize novel view, but also perform various downstream tasks, such as 1) HDR exposure editing, 2) Image deblurring.
[…]
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
1. Introduction
@@
Photo-realistic Novel View Synthesis (NVS) is a vital area in computer vision, enabling applications such as virtual reality/augmented reality (VR/AR), autonomous driving, and embodied AI by providing immersive visual experiences and advanced perception capabilities. Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), have gained prominence in NVS due to their high-quality 3D scene rendering, inspiring a wealth of subsequent research (Müller et al., 2022; Zhao et al., 2024; Wang et al., 2023; Huang et al., 2022).
However, most NVS methods rely on well-exposed, low dynamic range (LDR) images with consistent lighting, struggling to handle high dynamic range (HDR) scenes where details in very bright or dark regions are lost. This limitation restricts the reconstruction of fine details in HDR environments, reducing NVS applicability in real-world scenarios. While 2D HDR content has been standardized and widely adopted (Hannuksela et al., 2015; ITU-R, 2018; Alakuijala et al., 2019), 3D HDR free-viewpoint content remains an emerging field with significant potential. Thus, reconstructing 3D HDR scenes is essential for improving visual fidelity and supporting downstream tasks.
@@
Current 3D HDR NVS methods can be divided into two categories. The first category, e.g. RawNeRF (Mildenhall et al., 2022) and LE3D (Jin et al., 2024) etc., takes in noisy RAW images, aiming to reconstruct noise-free 3D HDR scenes. The second category, represented by (Huang et al., 2022; Cai et al., 2024; Jun-Seong et al., 2022; Wang et al., 2024b; Wu et al., 2024), draws inspiration from HDR imaging (HDRI), using multi-exposure LDR images as inputs to reconstruct the 3D HDR scene while learning camera response function (CRF). However, the strict inputs limit their flexibility and broader applications. The challenges include:
  • Data acquisition of RAW images and accurate exposure time is usually expensive due to the use of professional equipment;
  • In low-light conditions, long exposure times increase the risk of motion blur from camera shake, reducing reconstruction quality;
  • The geometric consistency will be compromised if given inaccurate camera pose initialization, as the camera poses are not being optimized.
Thus, a key challenge is reducing the cost of data acquisition, enabling high-quality and robust 3D HDR scene reconstruction with consumer-grade devices.
@@
Most modern consumer-grade cameras use auto-exposure (AE) during video recording, automatically adjusting exposure time based on ambient lighting. This expands the captured dynamic range in the video, making it possible to reconstruct 3D HDR scenes. However, naively applying these videos to existing HDR 3D reconstruction methods presents several challenges:
  • Although some cameras support providing exposure time in images stored in EXIF metadata, it is often impossible to obtain the exposure time for each frame in a video format;
  • AE can cause inconsistencies in brightness between frames, leading to pose estimation errors in structure from motion (SfM) frameworks;
  • Videos suffer from severe motion blur due to the camera movement during longer exposure times.
Addressing these issues requires modeling camera motion during exposure, rather than assuming a static camera as in prior work. Motion blur and brightness variations both depend on exposure time: longer exposures increase blur and brightness. Previous studies (Huang et al., 2022; Wang et al., 2023) have decoupled these effects into camera movement and irradiance accumulation. Building on this, we introduce a continuous-time trajectory on $\mathbb{SE}(3)$ to represent video camera motions, enabling joint optimization of poses and exposure times. We also design a CRF module to convert accumulated irradiance (HDR) into LDR image brightness. By integrating camera motion into a unified differentiable rendering framework, we jointly optimize the trajectory, CRF, and exposure times.
@@
  • Casual3DHDR: A one-stage pipeline build upon a unified imaging model that leverages a continuous-time camera trajectory to jointly optimize poses, CRF, exposure times, and 3D HDR representation, enabling low-cost reconstruction from casually-captured auto-exposure (AE) videos.
  • A dataset of synthetic and real-world videos with significant brightness variations and motion blur, supporting further research.
@@
  • Extensive experiments demonstrating SOTA performance in reconstructing high-quality HDR scenes from casual videos.
2. Related Works
2.1. High Dynamic Range Imaging
@@
High Dynamic Range Imaging (HDRI) enhances luminosity beyond standard digital imaging by merging multi-exposure LDR images. In video capture, alternating long and short exposures achieves similar effects. Recently, deep learning approaches have treated HDRI as an image domian translation task, designing networks to convert LDR to HDR images. However, camera motion during exposure time often lead to ghosting artifacts. To address this, (Gryaditskaya et al., 2015) proposed an adaptive metering algorithm to adjust exposure and reduce motion artifacts, while other methods use spatial attention to mitigate motion blur. With the advent of 3D scene representations such as NeRF and 3DGS, methods like (Huang et al., 2022; Jun-Seong et al., 2022; Cai et al., 2024; Huang et al., 2024; Wang et al., 2024b) have emerged to reconstruct 3D HDR scenes and calibrate CRFs simultaneously. While these methods are effective, they often rely on precise exposure times and struggle with motion blur, highlighting the need for improved robustness.
2.2. Image deblurring
@@
Image deblurring aims to restore sharp images from blurred ones and can be categorized into three types. The first type uses hand-crafted priors, like total variation and heavy-tailed gradient priors, to solve for the blur kernel (Krishnan and Fergus, 2009; Cho and Lee, 2009), but struggles with different kernels producing similar effects. The second type, deep learning-based, achieves end-to-end restoration using large datasets (Zamir et al., 2021; Tsai et al., 2022). The third type leverages multi-view blurred images to reconstruct the 3D scene and deblur the images with geometric constraints. Works like (Ma et al., 2021; Peng and Chellappa, 2022; Lee et al., 2023a) jointly learn the blur kernel and radiance field, while BAD-NeRF (Wang et al., 2023) introduces a physical motion blur model that jointly optimizes the radiance field and camera trajectories. Following BAD-NeRF (Wang et al., 2023), emerging works (Lee et al., 2023b; Li et al., 2024c; Lee et al., 2024b, a; Sun et al., 2024; Chen and Liu, 2024; Oh et al., 2024; Zhao et al., 2024; Yu et al., 2024; Li et al., 2024b; Qi et al., 2024; Tang et al., 2024) have demonstrated the effectiveness of using $\mathbb{SE}(3)$ trajectory representations for joint 3D reconstruction and blur image recovery.
2.3. Robust Novel View Synthesis
@@
Novel view synthesis aims to generate images from arbitrary viewpoints using input images with known poses. Neural Radiance Fields (NeRF)(Martin-Brualla et al., 2021) and 3D Gaussian Splatting (3DGS)(Kerbl et al., 2023) advanced this field by reconstructing 3D scene as radiance field and achieve great success.
[…]
3. Methods
@@
In this section, we will provide a detailed explanation of our proposed Casual3DHDR, which takes an AE video as input. We first use continuous-time trajectory to represent the camera motion during video recording rather than addressing each camera pose individually, so that the camera pose and exposure time can be jointly optimized through bundle adjustment. Subsequently, we design a camera response function (CRF) module which can convert the accumulation of irradiance (HDR image) over the exposure time to pixel value (LDR image). Due to the fact that the exposure time is the connection between the continuous-time trajectory and CRF module, we can treat them as a unified differentiable imaging model thus the trajectory, CRF and exposure time can be jointly optimized and constrained together.
3.1. Preliminary: 3D Gaussian Splatting
@@
(1) $\mathbf{G}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{\top}\mathbf{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})}$
(2) $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T},\mathbf{\Sigma^{\prime}}=\mathbf{J}\mathbf{R}_{c}\mathbf{\Sigma}\mathbf{R}_{c}^{T}\mathbf{J}^{T},$
@@
(3) $\mathbf{C}(x,y,\mathbf{P}_{i})=\sum_{i=1}^{N}\mathbf{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$
(4) $\alpha_{i}=\mathbf{o}_{i}\cdot\exp(-\sigma_{i}),\quad\sigma_{i}=\frac{1}{2}\Delta_{i}^{T}\mathbf{\Sigma^{\prime}}^{-1}\Delta_{i},$
@@
Figure 2. The pipeline of Casual3DHDR. Given a casually-captured video with AE, camera motion blur, and significant exposure time changes, we train 3DGS to reconstruct an HDR scene. We design a unified model based on the physical image formation process, integrating camera motion blur and exposure-induced brightness variations. This allows for the joint estimation of camera motion, exposure time, and camera response curve while reconstructing the HDR scene. After training, our method can sharpen the train images and render HDR images
3.2. Continuous Trajectory Representation
@@
Cumulative $\mathbb{SE}(3)$ B-splines are widely used in robotics for continuous-time trajectory representation, particularly in state estimation, sensor fusion, and path planning (Furgale et al., 2012; Lovegrove et al., 2013; Bry et al., 2015; Rehder et al., 2016; Mueggler et al., 2018; Geneva et al., 2020). These splines offer desirable properties, including $C^{2}$ continuity, locality, and the convex hull property, which effectively incorporate gradient information and dynamic constraints, enabling rapid convergence to smooth, feasible trajectories (Zhou et al., 2019). Moreover, $\mathbb{SE}(3)$ B-splines allow the computation of pose, velocity, and acceleration at any timestamp along the trajectory.
@@
Existing multi-view deblurring methods, such as BAD-NeRF (Wang et al., 2023), model camera motion by estimating short splines for each frame independently, limiting their ability to leverage cross-frame motion constraints and priors in continuous video inputs. Other approaches (Wang et al., 2021; Li et al., 2022; Sun et al., 2024; Li et al., 2024a; Lin et al., 2024; Shih et al., 2024; Wang et al., 2024c) employ basis functions to regularize continuous-time deformations for dynamic scene reconstruction. However, reconstructing scenes from casually-captured videos using a continuous-time camera trajectory representation remains unexplored. To address this, we propose modeling camera motion across an entire video using a cumulative $\mathbb{SE}(3)$ B-spline trajectory, enabling robust and consistent trajectory estimation.
@@
Following (Lovegrove et al., 2013), given a series of temporally uniformly distributed control knots, the pose $\mathbf{P}(t)$ at a given timestamp $t$ can be interpolated with 4 adjacent control knots, denoted as $\mathbf{T}_{\mathrm{0}}$, $\mathbf{T}_{1}$, $\mathbf{T}_{2}$ , $\mathbf{T}_{3}\in\mathbb{SE}(3)$:
(5) $\mathbf{P}(t)=\mathbf{T}_{0}\cdot\prod_{j=0}^{2}\mathrm{exp}(\tilde{\mathbf{B}}(u)_{j+1}\cdot\bm{\Omega}_{j}),$
(6) $\tilde{\mathbf{B}}(u)=\mathbf{C}\begin{bmatrix}1\\ u\\ u^{2}\\ u^{3}\end{bmatrix},\quad\mathbf{C}=\frac{1}{6}\begin{bmatrix}6&0&0&0\\ 5&3&-3&1\\ 1&3&3&-2\\ 0&0&0&1\end{bmatrix}$.
@@
where $\tau$ represents the spline sampling interval, $u=\frac{t}{\tau}$, and $u$ lies within the interval $[0,1)$; $\tilde{\mathbf{B}}(u)_{j+1}$ denotes the $(j+1)^{th}$ element of the vector $\tilde{\mathbf{B}}(u)$, $\bm{\Omega}_{j}=\mathrm{log}(\mathbf{T}_{j}^{-1}\cdot\mathbf{T}_{j+1})$, based on the (Qin, 1998).
Figure 3. Qualitative comparison on the Girls-vicon sequence of the Realsense dataset in terms of NVS.
Figure 4. Qualitative comparison on the Building sequence of the Smartphone dataset in terms of NVS.
3.3. Physical Image Formation Model